You have $35,000 to invest in a stock portfolio. Your choices are Stock X with an expected return of 14 percent and Stock Y with an expected return of 8 percent.

  

Required:
(a)

If your goal is to create a portfolio with an expected return of 11.1 percent, how much money will you invest in Stock X?

(Click to select)$64,750$18,987$17,179$18,806$18,083

  

(b)

If your goal is to create a portfolio with an expected return of 11.1 percent, how much money will you invest in Stock Y?

(Click to select)$17,594$16,071$16,917$17,763$16,240
Answer #1

Let investment in X=$x

Hence investment in Y=(35000-x)

Portfolio return=Respective return*Respective investment weight

11.1=(x/35000*14)+(35000-x)/35000*8

(11.1*35000)=14x+280,000-8x

388500=14x+280,000-8x

x=(388500-280,000)/(14-8)

=$18083(Approx)=investment in X

Hence investment in Y=(35000-18083)

=$16917(Approx).